In the reaction CH4(g)+2O2(g)-> CO2(g)+ 2H2O(l)
a. How many milligrams of CH4 react with 24.7mg of O2?
b.How many milligrams of CO2 are produced from 24.7mg of O2?
1
Expert's answer
2020-03-25T12:47:03-0400

(a)

From the reaction we can see that the ratio of stoichiometric coefficient of methane to oxygen is in ratio 1:2

Which means that for every part of oxygen half part methane is consumed

Moles of oxygen =24.7/32 millimoles

Moles of methane needed will be half of this which will be equal to =24.7/64 millimoles

Converting this into mass =(24.7/64)*16=6.2 mg


(b)

From the reaction we can see that the ratio of stoichiometric coefficient of oxygen to carbondioxide is in ratio 2:1

which means that for every mole of oxygen half mole of carbondioxide is produced

moles of oxygen ="\frac{24.7}{32}millimoles"


moles of carbondioxide produced = "\frac{24.7}{32\times2}millimoles"

converting this into mass= "\frac{24.7}{32\times2}\times44\ mg=16.98 mg"